A Bibliography of Publications in Biometrics: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 June 2020 Version 1.00

Title word cross-reference

#33109 [181].

2 [71]. 2^{3} [99]. 2 × 2 [127]. 3^{2} [99]. AOV [144]. K [51]. S [66]. ∑(O − E)^{2}/E [43]. t [86].

-based [178]. -method [66]. -sample [51].

29 [180].

30 [170, 179]. 31 [172]. 32 [116, 171, 146, 148, 149, 182, 147]. 33 [181]. 35 [215].

58 [181].

8 [185]. 81a [215].

A.O.V. [77]. abbreviated [6]. absence [95]. absorption [212, 139]. abundance [194]. Acceptance [13]. acute [32]. Advances [30]. age [121, 53]. age-dependent [121, 53]. aggregation [67]. agreement [95]. allele [162, 181, 16, 80]. allocation [165]. allowing [121]. Alternative [184, 151]. Analogies [110]. analyse [46]. analyses [24]. Analysis [207, 125, 113, 10, 82, 12, 108, 89, 192, 100, 126, 60, 81, 50, 46, 14, 93, 177, 211, 94, 66]. analytical [77]. Andrews [30]. angular [97]. ANOVA [75, 141, 55]. AOV [171]. Application [66, 160, 32, 201, 40, 180]. applications [175, 92]. approach [151]. Approaches [62]. approximate [74]. approximation [43, 201]. Approximations [39, 204]. arbitrary [64, 170]. arising [7]. ascertainment [161]. aspects [173]. Assumptions [161]. Asymptotic [20, 60]. Atlanta [32]. attribute [174]. augmented [78, 87]. available [120, 158, 159, 182].

balanced [138]. band [121]. band-return [121]. bands [34]. based [31, 178, 97, 126]. basis [19, 211]. Bayes [20, 41]. Bayesian [154, 21, 151]. behavior [20]. between [56, 95, 64, 170, 97, 110]. bias [39, 148, 149, 111, 112]. biased [175]. BIBD [42]. bibliography [117]. Bickel [30]. binary [174]. binomial [74, 2, 19, 145]. biometric [92]. Biometrics [116, 181, 170, 215, 171, 146, 148, 149, 185, 172, 179, 182, 180, 147]. birth [212, 17, 139]. birth-death [212, 139]. bivariate [102, 24]. Bliss [210]. block [57, 138, 133, 42]. blocks [100]. blood [103]. both [73, 65]. Bradley [116, 168, 151]. branching [160, 102, 53].

C [13]. Calculation [6, 71]. canonical [36]. capture [39]. carcass [22]. caries [203]. case [208, 64, 170]. case-control [208]. cases [32]. categorical [167]. category [129]. censored [215, 197, 183]. censorship [196]. censuses [1, 185]. chain [74]. chain-binomial [74]. chains [214]. Chapman [185]. characterization [28]. Chester [210]. chi [106]. chi-square [106]. choice [138]. class [65, 7, 142, 94]. classification [198, 31, 25, 79, 132]. classifications [72]. classified [73, 113]. classifying [20]. clinical [165]. clustering [32]. coefficients [3, 91]. column [87]. combinations [124]. Combining [129]. common [213]. Comparative [154, 22]. Comparing [196, 165]. comparison [140, 109, 144, 171]. comparisons [116, 115, 75, 117, 86, 77, 151, 153]. compartment [146, 123]. compartmental [128]. Competing [8, 200, 82]. complete [57, 100, 133, 42]. completely [73]. component [140, 10, 178, 11]. components [155, 37, 71, 33, 167, 176, 40, 180]. Composite [119]. composition [22]. comprehensive [211]. Computation [55, 98, 3]. Computing [155]. conditional [8, 81]. Confidence [124, 81]. Conservatism [43]. constants [146, 123]. constraints [83]. contingency [84, 198, 199, 73, 113, 5, 106, 110]. continuous [56, 214]. continuous-time [214]. contrasts [163, 153]. control [45, 208]. Cook [44]. Correction [215]. Corrections [181, 170, 171, 146, 148, 149, 172, 179, 182, 180, 147]. correlated [143, 169, 142, 133, 24]. Correlation [27, 56, 36, 65, 204, 91]. correlations [202]. Cost [79]. Covariance [29, 176, 40, 180, 110, 94]. Covariances [64, 170, 55]. covariates [200, 190]. Cox [189, 187]. criterion [134]. critique [82]. Cross [138, 73]. cross-classified [73]. Cross-validatory [138]. culling [40, 180]. curve [196, 166]. curves [34, 165].

D [30, 185]. damped [52]. data [154, 75, 121, 122, 73, 113, 86, 103, 196, 215, 197, 173, 174, 167, 209, 43, 202, 158, 159, 183, 177, 159]. death [212, 17, 139]. decision [41]. Demographic [90]. densities [19, 80]. density [127, 59]. density-dependent [127]. dental [203]. dependent [20, 121, 127, 128, 53, 91]. derivation [36]. description [32]. Design [120, 182, 21, 101]. designs [57, 78, 87, 138, 133, 42, 156, 24]. detection [67]. determine [215, 197]. determined [127]. Diagnosis [90]. diagnostic [136, 203]. dichotomiques [46]. dichotomous [46]. different [92, 161]. differential [103]. diffusion [80]. dimensional [199, 73]. direct [58]. disadvantageous [4]. discrete [160, 56, 41]. Discriminant [192, 126, 60, 81, 134]. discs [54]. Discussion [116]. disease [67]. disequilibrium [64, 170]. distance [97]. distribution [21, 2, 53, 215, 197, 54, 38, 15, 18]. Distributions [90, 175]. diversity [152]. domains [195]. Doolittle [6]. dose [68]. Double [22, 9]. double-stage [9]. Dunnett [153]. Dynamics [16].

ecological [70]. ecology [194]. effects [168, 124]. Efficient [177]. elements [127]. emphasizing [153]. Empirical [90]. epidemics [160, 74]. epidemiological [7]. epistasis [64, 170]. epistemological [173]. equal [148, 149, 111, 112]. Equality [91]. equations [36, 176]. equivalence [208, 99]. Erratum [185]. error [60, 203]. estimate [96, 172, 120, 182]. estimated [20]. Estimates [30, 129, 155]. Estimating [53, 146, 123, 19, 71, 60]. Estimation [98, 160, 163, 56, 65, 195, 9, 21, 178, 138, 33, 107, 92, 203, 176, 11, 40, 158, 159, 180]. estimator [206, 105, 37, 213]. estimators [140, 131, 77, 59, 141, 35]. evaluating [136]. evaluation [79]. exact [114, 35]. examples [149, 111]. expectations [55]. expected [16, 6]. experiment [65, 51]. experiments [21, 10, 120, 182, 177, 24]. exponential [105, 52, 165]. Extended [133, 42, 57]. extending [168]. extension [153].

F [13, 30]. factor [211]. factor-analysis [211]. factorial [10]. familial [202]. families [175]. family [152]. favorable [162, 181]. fertilizing [186]. finite [16, 212, 88, 139, 214]. firing [63]. fisheries [69]. Fisherman [118]. fit [183]. fixation [162, 181, 16]. Fixed [89, 124, 148, 112]. folded [2]. form [94]. formation [135]. formulations [184]. fractional [10]. French [75, 47, 46]. frequencies [163, 97]. frequency [72]. function [134]. functions [27, 109].

G [185]. gain [132]. gamma [154]. gamma-modeled [154]. gene [97]. general [137, 128, 28]. generalization [206, 136, 153, 147, 130]. generalized [99]. generated [42]. generation [53]. generation-time [53]. generations [159]. Genetic [26, 191, 155, 97, 143, 169, 50, 193, 48]. genetics [161]. George [76, 85]. Georgia [32]. Gertrude [189, 187]. given [26]. glimpses [85]. goal [107]. goodness [183]. graph [3]. Group [107, 103, 145]. grouped [157]. grouping [207]. growth [98, 21, 34, 127, 52, 101, 166]. Grubbs [35].

H [30]. Haenszel [208, 213]. Hampei [30]. haploid [16]. Hardy [114]. having [198, 94]. Hayne [206]. Hazard [190]. health [62]. Henry [164]. heritability [65, 120, 158, 159, 182]. heron [209]. heterogeneous [75, 177]. Higher [90]. Higher-Order [90]. homogeneity [99, 25]. homogeneous [212, 139, 214]. household [67]. Huber [30]. human [175, 161]. humans [191]. Hybrid [23]. hypotheses [208, 35].

II [148, 169, 112, 159]. Illustrated [90]. implications [143, 169]. improved [162, 181, 105, 37, 104]. inbreeding [3]. incidence [43, 203]. Incomplete [5, 138, 100]. incompletely [113]. incorporate [168]. independent [84, 24]. Index [83]. inequality [48]. inference [154]. Inferences [38, 122]. information [26]. integral [201]. inter [138]. inter- [138]. interaction [144, 171]. interclass [202]. internal [205]. interpretation [173]. intervals [86, 124, 81]. intra [65, 94]. intra-class [65, 94]. intrablock [138]. invariance [71]. Invariant [141, 131, 25]. Inverse [1, 19, 185, 18, 58]. involving [24]. isotonic [49]. iterative [155]. Ittner [210].

J [13, 30]. jackknife [68]. Jolly [39]. Jr. [164]. judges [95].

kidney [49]. kinship [3]. knowledge [9]. known [196]. Kolmogorov [136].

large [213]. lattice [156]. Laurence [164]. law [114, 52]. least [155, 205]. Leslie [98, 127, 69]. leukemia [32]. life [196, 104, 207, 157]. like [70]. likelihood [183]. line [122, 206, 59, 61, 179]. linear [137, 72, 124, 108, 134, 92, 28]. linkage [64, 170]. locally [2]. Location [30]. Log [72, 125, 163]. log-frequencies [163]. Log-linear [72]. Logical [173]. logistic [38]. logit [208, 12]. logrank [43]. loss [16]. Lotka [98]. Lucas [164]. luck [118].

M [189]. Mantel [208, 213]. marginal [199, 99]. Markov [214]. Mary [187]. matching [148, 149, 111, 112]. Mathematical [122, 45]. matrix [127, 69, 147, 130]. Maximally [131]. Maximally-invariant [131]. Maximum [132]. Maximums [148, 112]. McNemar [99]. mean [9, 6, 55]. measures [152]. Measuring [95, 32]. metastases [135]. method [125, 117, 178, 104, 59, 61, 179, 205, 66, 71]. methodology [150]. methods [6, 148, 149, 111, 112]. metropolitan [32]. minimum [106]. misallocation [81]. missing [23]. misspecification [28]. mixed [129, 184, 77, 89, 176, 58]. model [198, 45, 168, 74, 128, 103, 39, 49, 184, 77, 144, 171, 146, 123, 141, 167, 151, 135, 166, 28, 176]. modeled [154]. Modelling [209]. Models [121, 98, 200, 191, 122, 194, 63, 82, 72, 69, 89, 80, 193, 190, 92, 161, 110]. modern [189]. moments [212]. morphometrics [166]. mortality [51]. most [2]. Mosteller [13]. MR [181, 215]. multi [80]. multi-allele [80]. multidimensional [46]. multidimensionnel [46]. Multifactorial [191, 50, 201]. multinomial [31, 7]. multinormal [142]. Multiple [153, 75, 1, 144, 171, 185]. multiplicative [110]. multitrait [176]. Multivariate [12, 148, 149, 111, 112, 14, 201, 15]. mutants [4].

Nassar [44]. nature [173]. nature-nurture [173]. near [68]. negative [19]. nested [145]. neuron [63]. neutral [80]. no [116, 181, 215, 171, 146, 148, 149, 172, 182, 147]. nonparametric [122]. Normal [204, 201, 15]. note [128, 106, 214, 156]. numerical [77]. nurture [173].

Obituary [76, 164]. observations [20, 23, 133, 120, 157, 182, 158]. observed [196]. occupancy [62, 147, 130]. odds [125, 213]. offspring [65, 120, 158, 182]. one [68, 136, 65, 141, 167]. one-sided [136]. one-way [141, 167]. optimum [145]. Order [143, 169, 90, 142, 168]. ordered [198, 72]. outlier [137].

P [13, 30]. pair [168, 103]. paired [116, 115, 117, 151]. paper [44]. parameter [105]. parameters [92, 38]. parametric [206, 100]. parent [65]. parents [120, 158, 182]. partially [73]. patient [177]. pattern [66]. percent [148, 149, 111, 112]. periodic [23]. permutational [7]. personal [85]. pest [45]. pioneer [189]. placed [54]. Poisson [21, 18]. polynomial [23]. pooling [150]. population [98, 129, 16, 127, 19, 88, 59, 41]. populations [16, 186, 175, 58, 177]. power [27]. powerful [2]. prediction [150, 50]. preliminary [126]. presence [95, 196, 203]. prior [9]. Probabilities [212, 8, 139]. probability [162, 181, 16, 81]. problem [101, 62, 147, 130]. problems [92, 15]. procedure [145, 41, 153, 79, 132]. procedures [184, 144, 171]. process [102, 53, 17]. processes [160, 212, 139]. Processing [47]. Professional [85]. programmes [143, 169]. Properties [59, 152]. proportionality [83]. proportions [129, 68].

Quadratic [33, 131, 37, 141]. quadric [52]. qualitative [50, 201]. quantitative [191]. Quasi [84]. Quasi-independent [84]. quasisymmetry [199].

R [30]. Ralph [116]. random [196, 124, 141]. Randomization [93]. randomly [215, 197, 54, 183]. Ranking [100]. rate [98, 146, 123, 190]. rates [21, 121, 60]. ratio [125, 213, 96, 172, 183]. reabsorption [49]. recapture [39, 51]. records [40, 180]. recruitment [51]. rectangular [156]. recurrence [50]. reducing [148, 149, 111, 112]. reduction [148, 112]. reflections [188]. Regression [125, 22, 137, 65, 108, 23, 28, 94]. regressions [150]. relationship [97]. relative [26]. relatives [64, 170]. Reliability [174]. réponses [46]. research [7, 62, 14]. residuals [94]. response [198, 109, 169]. responses [68, 46, 24]. results [102]. retrospective [125]. return [121]. Review [13, 30, 82]. revisited [68]. risk [82]. risks [200, 8, 50]. Robust [30]. Rogers [30]. root [6]. row [87]. row-column [87]. rules [20, 41]. run [70]. run-like [70].

saline [49]. Sample [31, 20, 1, 136, 213, 51, 148, 112, 185]. Sample-based [31]. samples [126]. Sampling [13, 152, 21, 206, 22, 19, 88, 175, 119, 58, 61, 96, 172, 179, 18]. Science [116, 115]. score [208]. Seber [39]. selected [153]. selecting [134]. Selection [186, 15, 29, 83, 143, 169, 108, 193, 110]. Selectively [4]. self [186]. self-fertilizing [186]. separation [97]. Sequential [165, 1, 185]. serial [204]. Sets [90, 124]. several [109]. sibs [65]. sided [136]. Significance [202]. simple [36, 74, 178]. Simultaneous [92, 176]. single [137, 63]. size [16, 175]. size-biased [175]. sizes [148, 112]. small [20, 126, 195]. small-sample [20]. Smirnov [136]. Snedecor [85]. solution [162, 181]. Some [188, 87, 77, 144, 171, 193, 149, 35, 24, 82, 102, 111]. spatial [66]. Spearman [105]. Spearman-type [105]. special [142]. square [6, 106]. squares [155, 6, 205, 55]. Stability [127]. stage [9, 79, 132]. standard [109]. statistic [136, 70, 183]. statistical [173, 32, 46]. statistics [189, 116, 115, 208, 143, 169, 204, 106, 142]. statistique [46]. stepwise [41]. stimulus [46]. Stochastic [200, 194, 63, 135, 128, 49, 146, 123]. structure [94]. studies [125, 22, 208, 69, 101, 166, 157]. subcategory [129]. subject [40, 180]. successive [96, 172]. suggested [86]. summary [75]. super [88]. super-population [88]. Survey [30, 63]. survival [154, 200, 121, 196, 165, 209, 43, 177]. symmetric [2]. symmetry [199]. Symposium [13]. synthesis [178]. synthetic [211].

table [196, 104]. Tables [90, 84, 198, 199, 73, 113, 72, 207, 5, 99, 106, 110]. tag [51]. tag-recapture [51]. technique [60]. temporally [24]. Terry [168, 151]. test [114, 2, 32, 126, 43, 25]. Testing [208, 137, 215, 197, 68, 145, 107, 202, 183]. Tests [51, 199, 208, 86, 136, 7, 99, 35]. theory [3, 211]. three [199]. three-dimensional [199]. time [20, 160, 16, 128, 97, 53, 214]. time-dependent [20, 128]. times [139]. Tolerance [34]. Topics [11]. traditional [154]. trains [63]. trait [95]. traitement [47]. traits [191, 50, 201]. transect [122, 206, 59, 61, 179]. transects [70]. transformations [97]. Transition [80]. treatment [157]. trends [14]. trials [165]. triangular [84]. tubules [49]. Tukey [13, 30]. tumor [43]. twin [103]. twin-pair [103]. Two [73, 90, 199, 95, 136, 184, 144, 171, 146, 123, 41, 165, 25, 159, 79, 132]. two- [199]. two-compartment [146, 123]. Two-dimensional [73]. two-sample [136]. two-stage [79, 132]. Two-Way [90, 184, 144, 171, 25]. type [105].

ultimate [162, 181]. Unbalanced [94, 158, 159]. unbiased [131, 60]. underlying [215, 197]. unequally [142]. unimodal [32]. Use [3, 96, 172]. used [82]. Using [68, 9, 215, 197, 41, 209].

validatory [138]. valued [17]. variable [16, 56]. variables [47, 143, 169, 108, 41, 134]. variance [140, 155, 37, 178, 213, 71, 89, 33, 167, 176, 11, 40, 180, 93]. variances [75, 25, 55]. variates [142]. vector [47, 17]. vector-valued [17]. vectorielles [47]. versus [58]. viability [103]. viewpoint [88].

W [85, 13, 30]. Way [90, 184, 144, 171, 141, 167, 25]. weather [209]. Weibull [177]. Weighted [175, 155]. weights [138]. Weinberg [114]. wildlife [59, 175]. Winsor [13]. withdrawals [207]. within [186, 168]. within-pair [168]. without [23]. work [14].

zero [68]. Zyskind [76].

References

Chapman:1952:IMS

[1] Douglas G. Chapman. Inverse, multiple and sequential sample censuses. Biometrics, 8(??):286–306, 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See erratum [185].

Gart:1970:LMP

[2] John J. Gart. A locally most powerful test for the symmetric folded binomial distribution. Biometrics, 26(??):129–138, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maruyama:1970:UGT

[3] Takeo Maruyama and Norikazu Yasuda. Use of graph theory in computation of inbreeding and kinship coefficients. Biometrics, 26(??):209–219, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thomson:1970:SDM

[4] Glenys J. Thomson. Selectively disadvantageous mutants. Biometrics, 26 (??):229–241, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1970:ICT

[5] Nathan Mantel. Incomplete contingency tables. Biometrics, 26(??):291– 304, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gaylor:1970:CEM

[6] D. W. Gaylor, H. L. Lucas, and R. L. Anderson. Calculation of expected mean squares by the abbreviated Doolittle and square root methods. Biometrics, 26(??):641–655, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1970:CPM

[7] Nathan Mantel and John C. Bailar III. A class of permutational and multinomial tests arising in epidemiological research. Biometrics, 26(??): 687–700, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiang:1970:CRC

[8] Chin Long Chiang. Competing risks and conditional probabilities. Biometrics, 26(??):767–776, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arnold:1970:DSE

[9] J. C. Arnold and H. A. Al-Bayyati. On double-stage estimation of the mean using prior knowledge. Biometrics, 26(??):787–800, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Das:1970:CAF

[10] M. N. Das and R. C. Jain. On component analysis of factorial and fractional factorial experiments. Biometrics, 26(??):823–833, 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Searle:1971:TVC

[11] S. R. Searle. Topics in variance component estimation. Biometrics, 27 (??):1–76, 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grizzle:1971:MLA

[12] James E. Grizzle. Multivariate logit analysis. Biometrics, 27(4):1057– 1062, 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

F:1971:RAS

[13] W. T. F. Review: Acceptance Sampling – A Symposium, by J. W. Tukey, F. Mosteller, and C. P. Winsor. Biometrics, 27(4):1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

http://www.jstor.org/stable/2528858.

Rao:1972:RTR

[14] C. Radhakrishna Rao. Recent trends of research work in multivariate analysis. Biometrics, 28(??):3–22, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1972:SPU

[15] Ying Y. Wang. Selection problems under multivariate normal distribution. Biometrics, 28(??):223–233, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:1972:DFP

[16] R. D. Cook and R. F. Nassar. Dynamics of finite populations. I. The expected time to fixation or loss and the probability of fixation of an allele in a haploid population of variable size. Biometrics, 28(??):373–384, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [44].

Shanbhag:1972:VVB

[17] D. N. Shanbhag. On a vector-valued birth and death process. Biometrics, 28(??):417–425, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weiler:1972:ISP

[18] H. Weiler. Inverse sampling of a Poisson distribution. Biometrics, 28(??): 959–970, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerrard:1972:IBS

[19] Douglas J. Gerrard and R. D. Cook. Inverse binomial sampling as a basis for estimating negative binomial population densities. Biometrics, 28(??): 971–980, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Azen:1972:ASS

[20] S. P. Azen and A. A. Afifi. Asymptotic and small-sample behavior of estimated Bayes rules for classifying time-dependent observations. Biometrics, 28(??):989–998, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Behnken:1972:BED

[21] Donald W. Behnken and Donald G. Watts. Bayesian estimation and design of experiments for growth rates when sampling from the Poisson distribution. Biometrics, 28(??):999–1009, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conniffe:1972:DSR

[22] D. Conniffe and M. A. Moran. Double sampling with regression in comparative studies of carcass composition. Biometrics, 28(??):1011–1023, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Housworth:1972:HPP

[23] William J. Housworth. Hybrid polynomial and periodic regression with and without missing observations. Biometrics, 28(??):1025–1042, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1972:SDA

[24] Sing Chou Wu, James S. Williams, and Paul W. Mielke, Jr. Some designs and analyses for temporally independent experiments involving correlated bivariate responses. Biometrics, 28(??):1043–1061, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shukla:1972:ITH

[25] G. K. Shukla. An invariant test for the homogeneity of variances in a two-way classification. Biometrics, 28(??):1063–1072, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jacquard:1972:GIG

[26] Albert Jacquard. Genetic information given by a relative. Biometrics, 28 (??):1101–1114, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Good:1972:CPF

[27] I. J. Good. Correlation for power functions. Biometrics, 28(??):1127–1129; correction, ibid. 29 (1973), 829, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenberg:1972:CMG

[28] Saul H. Rosenberg and Paul S. Levy. A characterization on misspecification in the general linear regression model. Biometrics, 28(??):1129–1133, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dempster:1972:CS

[29] A. P. Dempster. Covariance selection. Biometrics, 28(1):157–175, 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robertson:1972:RRE

[30] C. A. Robertson. Review: Robust Estimates of Location: Survey and Advances, by D. R. Andrews, P. J. Bickel, F. R. Hampei, P. J. Huber, W. H. Rogers, and J. W. Tukey. Biometrics, 28(4):1144–1145, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glick:1973:SBM

[31] Ned Glick. Sample-based multinomial classification. Biometrics, 29(??): 241–256, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Larsen:1973:STM

[32] Richard J. Larsen, Charles L. Holmes, and Clark W. Heath, Jr. A statistical test for measuring unimodal clustering: a description of the test and of its application to cases of acute leukemia in metropolitan Atlanta, Georgia. Biometrics, 29(??):301–309, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LaMotte:1973:QEV

[33] L. R. LaMotte. Quadratic estimation of variance components. Biometrics, 29(??):311–330, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bowden:1973:TBG

[34] David C. Bowden and R. Kirk Steinhorst. Tolerance bands for growth curves. Biometrics, 29(??):361–371, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shukla:1973:SET

[35] G. K. Shukla. Some exact tests of hypotheses about Grubbs’s estimators. Biometrics, 29(??):373–377, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cramer:1973:SDC

[36] Elliot M. Cramer. A simple derivation of the canonical correlation equations. Biometrics, 29(??):379–380, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haile:1973:IQE

[37] Alemayehu Haile and J. T. Webster. An improved quadratic estimator for variance components. Biometrics, 29(??):408–413, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schafer:1973:IPL

[38] R. E. Schafer and T. S. Sheffield. Inferences on the parameters of the logistic distribution. Biometrics, 29(??):449–455, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:1973:ABJ

[39] Richard O. Gilbert. Approximations of the bias in the Jolly–Seber capture–recapture model. Biometrics, 29(??):501–526, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1973:EVC

[40] R. Thompson. The estimation of variance and covariance components with an application when records are subject to culling. Biometrics, 29 (??):527–550, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [180].

Lachin:1973:SPT

[41] John M. Lachin. On a stepwise procedure for two population Bayes decision rules using discrete variables. Biometrics, 29(??):551–564, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trail:1973:ECB

[42] Stanley M. Trail and David L. Weeks. Extended complete block designs generated by BIBD. Biometrics, 29(??):565–578, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peto:1973:CAL

[43] Richard Peto and M. C. Pike. Conservatism of the approximation ∑(O − E)^{2}/E in the logrank test for survival data or tumor incidence data. Biometrics, 29(??):579–584, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Watterson:1973:RPC

[44] G. A. Watterson. On a recent paper by Cook and Nassar. Biometrics, 29(??):595–600, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [16].

Chatterjee:1973:MMP

[45] Samprit Chatterjee. A mathematical model for pest control. Biometrics, 29(??):727–734, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meste:1973:ASR

[46] M. Meste. Sur l’analyse statistique des réponses dichotomiques à un stimulus multidimensionnel. (French) [On the statistical analysis of dichotomous responses to a multidimensional stimulus]. Biometrics, 29(??):735–749, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Escoufier:1973:TVV

[47] Yves Escoufier. Le traitement des variables vectorielles. (French) [Processing vector variables]. Biometrics, 29(??):751–760, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seneta:1973:GI

[48] E. Seneta. On a genetic inequality. Biometrics, 29(??):810–814, 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hardy:1974:SMR

[49] Robert J. Hardy and Edward J. Weinman. A stochastic model for the reabsorption of isotonic saline by the tubules in the kidney. Biometrics, 30(??):33–40, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mendell:1974:MQT

[50] Nancy Role Mendell and Robert C. Elston. Multifactorial qualitative traits: genetic analysis and prediction of recurrence risks. Biometrics, 30 (??):41–57, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pollock:1974:TMR

[51] K. H. Pollock, D. L. Solomon, and D. S. Robson. Tests for mortality and recruitment in a K-sample tag-recapture experiment. Biometrics, 30 (??):77–87, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fletcher:1974:QLD

[52] R. Ian Fletcher. The quadric law of damped exponential growth. Biometrics, 30(??):111–124, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoel:1974:EGT

[53] David G. Hoel and Kenny S. Crump. Estimating the generation-time distribution of an age-dependent branching process. Biometrics, 30(??): 125–135, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mohn:1974:DRP

[54] Erik Mohn and Per Stavem. On the distribution of randomly placed discs. Biometrics, 30(??):137–156, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Speed:1974:CEV

[55] F. M. Speed and R. R. Hocking. Computation of expectations, variances and covariances of ANOVA mean squares. Biometrics, 30(??):157–169, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1974:ECB

[56] N. R. Cox. Estimation of the correlation between a continuous and a discrete variable. Biometrics, 30(??):171–178, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornell:1974:MEC

[57] John A. Cornell. More on extended complete block designs. Biometrics, 30(??):179–186, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scheaffer:1974:DVI

[58] R. L. Scheaffer. On direct versus inverse sampling from mixed populations. Biometrics, 30(??):187–198, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kovner:1974:PEW

[59] J. L. Kovner and S. A. Patil. Properties of estimators of wildlife population density for the line transect method. Biometrics, 30(??):225–230, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLachlan:1974:AUT

[60] G. J. McLachlan. An asymptotic unbiased technique for estimating the error rates in discriminant analysis. Biometrics, 30(??):239–249, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sen:1974:LTS

[61] A. R. Sen, J. Tourigny, and G. E. J. Smith. On the line transect sampling method. Biometrics, 30(??):329–340, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [179].

Mantel:1974:AHR

[62] Nathan Mantel. Approaches to a health research occupancy problem. Biometrics, 30(??):355–362, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fienberg:1974:SMS

[63] Stephen E. Fienberg. Stochastic models for single neuron firing trains: a survey. Biometrics, 30(??):399–427, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gallais:1974:CBA

[64] A. Gallais. Covariances between arbitrary relatives with linkage and epistasis in the case of linkage disequilibrium. Biometrics, 30(??):429–446, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [170].

Hill:1974:EHB

[65] W. G. Hill and F. W. Nicholas. Estimation of heritability by both regression of offspring on parent and intra-class correlation of sibs in one experiment. Biometrics, 30(??):447–468, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zahl:1974:AMA

[66] Samuel Zahl. Application of the S-method to the analysis of spatial pattern. Biometrics, 30(??):513–524, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walter:1974:DHA

[67] S. D. Walter. On the detection of household aggregation of disease. Biometrics, 30(??):525–538, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frawley:1974:UJT

[68] W. H. Frawley. Using the jackknife in testing dose responses in proportions near zero or one—revisited. Biometrics, 30(??):539–545, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jensen:1974:LMM

[69] A. L. Jensen. Leslie matrix models for fisheries studies. Biometrics, 30 (??):547–551, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Knight:1974:RLS

[70] William Knight. A run-like statistic for ecological transects. Biometrics, 30(??):553–555, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henderson:1974:ICM

[71] C. R. Henderson, S. R. Searle, and L. R. Schaeffer. The invariance and calculation of Method 2 for estimating variance components. Biometrics, 30(??):583–588, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haberman:1974:LLM

[72] Shelby J. Haberman. Log-linear models for frequency tables with ordered classifications. Biometrics, 30(??):589–600, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1974:TDC

[73] Tar Chen and Stephen E. Fienberg. Two-dimensional contingency tables with both completely and partially cross-classified data. Biometrics, 30 (??):629–642, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dayananda:1974:ACB

[74] P. W. A. Dayananda. An approximate chain-binomial model for simple epidemics. Biometrics, 30(??):705–708, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:1974:AMC

[75] Morton B. Brown and Alan B. Forsythe. The ANOVA and multiple comparisons for data with heterogeneous variances. (French summary). Biometrics, 30(??):719–724, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1974:OGZ

[76] Anonymous. Obituary: George Zyskind (1929–1974). Biometrics, 30 (??):725, 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hocking:1975:SAN

[77] R. R. Hocking and M. H. Kutner. Some analytical and numerical comparisons of estimators for the mixed A.O.V. model. Biometrics, 31(??):19–28, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1975:AD

[78] W. T. Federer and D. Raghavarao. On augmented designs. Biometrics, 31 (??):29–35, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zielezny:1975:CET

[79] Maria Zielezny and Olive Jean Dunn. Cost evaluation of a two-stage classification procedure. Biometrics, 31(??):37–47, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Littler:1975:TDN

[80] R. A. Littler and E. D. Fackerell. Transition densities for neutral multiallele diffusion models. Biometrics, 31(??):117–123, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLachlan:1975:CIC

[81] G. J. McLachlan. Confidence intervals for the conditional probability of misallocation in discriminant analysis. Biometrics, 31(??):161–167, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gail:1975:RCS

[82] Mitchell Gail. A review and critique of some models used in competing risk analysis. Biometrics, 31(??):209–222, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harville:1975:ISP

[83] D. A. Harville. Index selection with proportionality constraints. Biometrics, 31(??):223–225, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altham:1975:QIT

[84] Patricia M. E. Altham. Quasi-independent triangular contingency tables. Biometrics, 31(??):233–238, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1975:PPG

[85] Gertrude M. Cox and Paul G. Homeyer. Professional and personal glimpses of George W. Snedecor. Biometrics, 31(??):265–301, 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duncan:1975:TIC

[86] David B. Duncan. t tests and intervals for comparisons suggested by the data. Biometrics, 31(??):339–359, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1975:SAR

[87] W. T. Federer, R. C. Nair, and D. Raghavarao. Some augmented rowcolumn designs. Biometrics, 31(??):361–373, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartley:1975:SPV

[88] H. O. Hartley and R. L. Sielken, Jr. A “super-population viewpoint” for finite population sampling. Biometrics, 31(??):411–422, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempthorne:1975:FMM

[89] Oscar Kempthorne. Fixed and mixed models in the analysis of variance. Biometrics, 31(??):473–486, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McNeil:1975:HOD

[90] D. R. McNeil and J. W. Tukey. Higher-order diagnosis of two-way tables, illustrated on two sets of demographic empirical distributions. Biometrics, 31(2):487–510, June 1975. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL http://www.jstor.org/stable/2529433.

Neill:1975:EDC

[91] John J. Neill and Olive Jean Dunn. Equality of dependent correlation coefficients. Biometrics, 31(??):531–543, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rao:1975:SEP

[92] C. Radhakrishna Rao. Simultaneous estimation of parameters in different linear models and applications to biometric problems. Biometrics, 31(??): 545–554, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

White:1975:RAV

[93] Robert F. White. Randomization and the analysis of variance. Biometrics, 31(??):555–571, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wiorkowski:1975:URA

[94] John J. Wiorkowski. Unbalanced regression analysis with residuals having a covariance structure of intra-class form. Biometrics, 31(3):611–618, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fleiss:1975:MAB

[95] Joseph L. Fleiss. Measuring agreement between two judges on the presence or absence of a trait. Biometrics, 31(3):651–659, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sen:1975:URE

[96] A. R. Sen, S. Sellers, and G. E. J. Smith. The use of a ratio estimate in successive sampling. Biometrics, 31(3):673–683, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [172].

Heuch:1975:RBS

[97] Ivar Heuch. The relationship between separation time and genetic distance based on angular transformations of gene frequencies. Biometrics, 31(3): 685–700, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:1975:ECG

[98] David H. Anderson. Estimation and computation of the growth rate in Leslie’s and Lotka’s population models. Biometrics, 31(3):701–718, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1975:EGM

[99] Nathan Mantel and Joseph L. Fleiss. The equivalence of the generalized McNemar tests for marginal homogeneity in 2^{3} and 3^{2} tables. Biometrics, 31(3):727–729, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Levitt:1975:RIC

[100] David J. Levitt. Ranking in incomplete or complete blocks: a parametric analysis. Biometrics, 31(3):745–748, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Machin:1975:DPG

[101] David Machin. On a design problem in growth studies. Biometrics, 31(3): 749–753, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghai:1975:SRB

[102] G. L. Ghai and E. Pollak. On some results for a bivariate branching process. Biometrics, 31(3):761–763, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fienberg:1975:DVM

[103] Stephen E. Fienberg and Michael I. Waller. A differential viability model for twin-pair blood group data. Biometrics, 31(3):785–790, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keyfitz:1975:ILT

[104] Nathan Keyfitz and James Frauenthal. An improved life table method. Biometrics, 31(4):889–899, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Church:1975:IST

[105] J. D. Church and E. Benton Cobb. An improved Spearman-type estimator for the exponential parameter. Biometrics, 31(4):913–920, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quade:1975:NMC

[106] Dana Quade and Ibrahim A. Salama. A note on minimum chi-square statistics in contingency tables. Biometrics, 31(4):953–956, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1975:GTG

[107] Nathan Mantel. Group testing with the goal of estimation. Biometrics, 31 (4):994–995, 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). With a reply by Sobel and Elashoff.

Hocking:1976:ASV

[108] R. R. Hocking. The analysis and selection of variables in linear regression. Biometrics, 32(1):1–49, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jensen:1976:CSR

[109] D. R. Jensen. The comparison of several response functions with a standard. Biometrics, 32(1):51–59, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wermuth:1976:ABM

[110] Nanny Wermuth. Analogies between multiplicative models in contingency tables and covariance selection. Biometrics, 32(1):95–108, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rubin:1976:MMMa

[111] Donald B. Rubin. Multivariate matching methods that are equal percent bias reducing. I. Some examples. Biometrics, 32(1):109–120, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [149].

Rubin:1976:MMMb

[112] Donald B. Rubin. Multivariate matching methods that are equal percent bias reducing. II. Maximums on bias reduction for fixed sample sizes. Biometrics, 32(1):121–132, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [148].

Chen:1976:ACT

[113] Tar Chen and Stephen E. Fienberg. The analysis of contingency tables with incompletely classified data. Biometrics, 32(1):133–144, 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapco:1976:ETH

[114] W. Chapco. An exact test of the Hardy–Weinberg law. Biometrics, 32(1): 183–189, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bradley:1976:SSP

[115] Ralph A. Bradley. Science, statistics, and paired comparisons. Biometrics, 32(2):213–232, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [116].

Beaver:1976:DSS

[116] R. J. Beaver, H. A. David, and Roger R. Davidson. Discussion: “Science, statistics and paired comparisons” (Biometrics 32 (1976), no. 2, 213–232) by Ralph A. Bradley. Biometrics, 32(2):233–239, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [115].

Davidson:1976:BMP

[117] Roger R. Davidson and Peter H. Farquhar. A bibliography on the method of paired comparisons. Biometrics, 32(2):241–252, 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1976:FL

[118] W. A. Thompson, Jr. Fisherman’s luck. Biometrics, 32(2):265–271, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rohde:1976:CS

[119] Charles A. Rohde. Composite sampling. Biometrics, 32(2):273–282, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1976:DEE

[120] Robin Thompson. Design of experiments to estimate heritability when observations are available on parents and offspring. Biometrics, 32(2): 283–304, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [182].

Brownie:1976:MAA

[121] C. Brownie and D. S. Robson. Models for allowing for age-dependent survival rates for band-return data. Biometrics, 32(2):305–323, 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burnham:1976:MMN

[122] K. P. Burnham and D. R. Anderson. Mathematical models for nonparametric inferences from line transect data. Biometrics, 32(2):325–336, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kodell:1976:ERC

[123] R. L. Kodell and J. H. Matis. Estimating the rate constants in a twocompartment stochastic model. Biometrics, 32(2):377–390, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections

Harville:1976:CIS

[124] David A. Harville. Confidence intervals and sets for linear combinations of fixed and random effects. Biometrics, 32(2):403–407, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Breslow:1976:RAL

[125] N. Breslow. Regression analysis of the log odds ratio: a method for retrospective studies. Biometrics, 32(2):409–416, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McDonald:1976:PTD

[126] Lyman L. McDonald, Victor W. Lowe, Robert K. Smidt, and Keren A. Meister. A preliminary test for discriminant analysis based on small samples. Biometrics, 32(2):417–422, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cooke:1976:SPG

[127] Dennis Cooke and Jesus Alberto Leon. Stability of population growth determined by 2 × 2 Leslie matrix with density-dependent elements. Biometrics, 32(2):435–442, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faddy:1976:NGT

[128] M. J. Faddy. A note on the general time-dependent stochastic compartmental model. Biometrics, 32(2):443–448, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:1976:CEC

[129] G. H. Brown. Combining estimates of category and subcategory proportions in a mixed population. Biometrics, 32(2):453–457, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walter:1976:GMO

[130] S. D. Walter. A generalization of a matrix occupancy problem. Biometrics, 32(2):471–475, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [147].

Goodnight:1976:MIQ

[131] James H. Goodnight. Maximally-invariant quadratic unbiased estimators. Biometrics, 32(2):477–480, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zielezny:1976:MGT

[132] Maria A. Zielezny. Maximum gain in the two-stage classification procedure. Biometrics, 32(2):481–484, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schreckengost:1976:ECB

[133] Jack F. Schreckengost and John A. Cornell. Extended complete block designs with correlated observations. Biometrics, 32(3):505–518, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLachlan:1976:CSV

[134] G. J. McLachlan. A criterion for selecting variables for the linear discriminant function. Biometrics, 32(3):529–534, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liotta:1976:SMM

[135] Lance A. Liotta, Gerald M. Saidel, and Jerome Kleinerman. Stochastic model of metastases formation. Biometrics, 32(3):535–550, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gail:1976:GOS

[136] Mitchell H. Gail and Sylvan B. Green. A generalization of the one-sided two-sample Kolmogorov–Smirnov statistic for evaluating diagnostic tests. Biometrics, 32(3):561–570, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ellenberg:1976:TSO

[137] Jonas H. Ellenberg. Testing for a single outlier from a general linear regression. Biometrics, 32(3):637–645, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jensen:1976:CVC

[138] Louis Jensen and Mervyn Stone. Cross-validatory choice of weights for inter- and intrablock estimation in balanced incomplete block designs. Biometrics, 32(3):677–681, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:1976:APA

[139] W. Y. Tan. On the absorption probabilities and absorption times of finite homogeneous birth-death processes. Biometrics, 32(4):745–752, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Corbeil:1976:CVC

[140] R. R. Corbeil and S. R. Searle. A comparison of variance component estimators. Biometrics, 32(4):779–791, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LaMotte:1976:IQE

[141] Lynn Roy LaMotte. Invariant quadratic estimators in the random, oneway ANOVA model. Biometrics, 32(4):793–804, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rawlings:1976:OSS

[142] J. O. Rawlings. Order statistics for a special class of unequally correlated multinormal variates. Biometrics, 32(4):875–887, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hill:1976:OSC

[143] William G. Hill. Order statistics of correlated variables and implications in genetic selection programmes. Biometrics, 32(4):889–902, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:1976:SNM

[144] Dallas E. Johnson. Some new multiple comparison procedures for the two-way AOV model with interaction. Biometrics, 32(4):929–934, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [171].

Hwang:1976:ONP

[145] F. K. Hwang. An optimum nested procedure in binomial group testing. Biometrics, 32(4):939–943, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kodell:1976:CER

[146] R. L. Kodell and J. H. Matis. Corrections to: “Estimating the rate constants in a two-compartment stochastic model” (Biometrics 32 (1976), no. 2, 377–390). Biometrics, 32(4):954, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [123].

Walter:1976:CGM

[147] S. D. Walter. Corrections to: “A generalization of a matrix occupancy problem” (Biometrics 32 (1976), no. 2, 471–475). Biometrics, 32(4):954, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [130].

Rubin:1976:CMMa

[148] D. B. Rubin. Corrections to: “Multivariate matching methods that are equal percent bias reducing. II. Maximums on bias reduction for fixed sample sizes” (Biometrics 32 (1976), no. 1, 121–132). Biometrics, 32 (4):955, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [112].

Rubin:1976:CMMb

[149] D. B. Rubin. Corrections to: “Multivariate matching methods that are equal percent bias reducing. I. Some examples” (Biometrics 32 (1976), no. 1, 109–120). Biometrics, 32(4):955, 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [111].

Johnson:1977:PMR

[150] J. P. Johnson, T. A. Bancroft, and Chien Pai Han. A pooling methodology for regressions in prediction. Biometrics, 33(1):57–67, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leonard:1977:ABA

[151] Tom Leonard. An alternative Bayesian approach to the Bradley–Terry model for paired comparisons. Biometrics, 33(1):121–132, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1977:SPF

[152] Woollcott Smith and J. Frederick Grassle. Sampling properties of a family of diversity measures. Biometrics, 33(2):283–292, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shaffer:1977:MCE

[153] Juliet Popper Shaffer. Multiple comparisons emphasizing selected contrasts: an extension and generalization of Dunnett’s procedure. Biometrics, 33(2):293–303, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartolucci:1977:CBT

[154] Alfred A. Bartolucci and James M. Dickey. Comparative Bayesian and traditional inference for gamma-modeled survival data. Biometrics, 33(2): 343–354, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornelius:1977:CIW

[155] P. L. Cornelius and J. Byars. Computing iterative weighted least squares estimates of genetic variance components. Biometrics, 33(2):375–382, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1977:NRL

[156] E. R. Williams. A note on rectangular lattice designs. Biometrics, 33(2): 410–414, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1977:TGO

[157] W. A. Thompson, Jr. On the treatment of grouped observations in life studies. Biometrics, 33(3):463–470, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1977:EHUa

[158] Robin Thompson. The estimation of heritability with unbalanced data. I. Observations available on parents and offspring. Biometrics, 33(3): 485–495, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1977:EHUb

[159] Robin Thompson. The estimation of heritability with unbalanced data. II. Data available on more than two generations. Biometrics, 33(3):497–504, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Becker:1977:EDT

[160] Niels Becker. Estimation for discrete time branching processes with application to epidemics. Biometrics, 33(3):515–522, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stene:1977:ADA

[161] Jon Stene. Assumptions for different ascertainment models in human genetics. Biometrics, 33(3):523–527, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cash:1977:ISU

[162] W. S. Cash. An improved solution for the ultimate probability of fixation of a favorable allele. Biometrics, 33(3):528–532, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [181].

Cook:1977:ECL

[163] G. W. Cook. Estimation of contrasts of log-frequencies. Biometrics, 33 (3):548–552, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1977:OHL

[164] Anonymous. Obituary: Henry Laurence Lucas, Jr. 1916–1977. Biometrics, 33(3):573, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Louis:1977:SAC

[165] Thomas A. Louis. Sequential allocation in clinical trials comparing two exponential survival curves. Biometrics, 33(4):627–634, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Madsen:1977:GCM

[166] K. S. Madsen. A growth curve model for studies in morphometrics. Biometrics, 33(4):659–669, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Landis:1977:OWC

[167] J. Richard Landis and Gary G. Koch. A one-way components of variance model for categorical data. Biometrics, 33(4):671–679, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davidson:1977:EBT

[168] Roger R. Davidson and Robert J. Beaver. On extending the Bradley– Terry model to incorporate within-pair order effects. Biometrics, 33(4): 693–702, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hill:1977:OSC

[169] William G. Hill. Order statistics of correlated variables and implications in genetic selection programmes. II. Response to selection. Biometrics, 33 (4):703–712, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gallais:1977:CCB

[170] A. Gallais. Corrections to: “Covariances between arbitrary relatives with linkage and epistasis, in the case of linkage disequilibrium” (Biometrics 30 (1974), 429–446). Biometrics, 33(4):766, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [64].

Johnson:1977:CSN

[171] Dallas E. Johnson. Corrections to: “Some new multiple comparison procedures for the two-way AOV model with interaction” (Biometrics 32 (1976), no. 4, 929–934). Biometrics, 33(4):766, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [144].

Sen:1977:CUR

[172] A. R. Sen, S. Sellers, and G. E. J. Smith. Corrections to: “The use of a ratio estimate in successive sampling” (Biometrics 31 (1975), no. 3, 673– 683). Biometrics, 33(4):767, 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [96].

Kempthorne:1978:LES

[173] Oscar Kempthorne. Logical, epistemological and statistical aspects of nature-nurture data interpretation. Biometrics, 34(1):1–23, 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krippendorff:1978:RBA

[174] Klaus Krippendorff. Reliability of binary attribute data. Biometrics, 34 (1):142–144, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). With a reply by Fleiss.

Patil:1978:WDS

[175] G. P. Patil and C. R. Rao. Weighted distributions and size-biased sampling with applications to wildlife populations and human families. Biometrics, 34(2):179–189, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schaeffer:1978:SEV

[176] L. R. Schaeffer, J. W. Wilton, and Robin Thompson. Simultaneous estimation of variance and covariance components from multitrait mixed model equations. Biometrics, 34(2):199–208, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1978:EAW

[177] J. S. Williams. Efficient analysis of Weibull survival data from experiments on heterogeneous patient populations. Biometrics, 34(2):209–222, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartley:1978:SSB

[178] H. O. Hartley, J. N. K. Rao, and L. R. LaMotte. A simple “synthesis”- based method of variance component estimation. Biometrics, 34(2):233– 242, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sen:1978:CLT

[179] A. R. Sen, J. Tourigny, and G. E. J. Smith. Corrections to: “On the line transect sampling method” (Biometrics 30 (1974), 329–340). Biometrics, 34(2):328–329, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [61].

Thompson:1978:CEV

[180] R. Thompson. Corrections to: “The estimation of variance and covariance components with an application when records are subject to culling” (Biometrics 29 (1973), 527–550). Biometrics, 34(2):328–329, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [40].

Cash:1978:CIS

[181] W. S. Cash. Corrections to: “An improved solution for the ultimate probability of fixation of a favorable allele” [Biometrics 33 (1977), no. 3, 528– 532; MR 58 #33109]. Biometrics, 34(2):329, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [162].

Thompson:1978:CDE

[182] R. Thompson. Corrections to: “Design of experiments to estimate heritability when observations are available on parents and offspring” (Biometrics 32 (1976), no. 2, 283–304). Biometrics, 34(2):329, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [120].

Turnbull:1978:LRS

[183] B. W. Turnbull and L. Weiss. A likelihood ratio statistic for testing goodness of fit with randomly censored data. Biometrics, 34(3):367–375, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harville:1978:AFP

[184] David A. Harville. Alternative formulations and procedures for the twoway mixed model. Biometrics, 34(3):441–453, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sandvik:1978:EIM

[185] Leiv Sandvik. Erratum: “Inverse, multiple and sequential sample censuses” (Biometrics 8 (1952), 286–306) by D. G. Chapman. Biometrics, 34(3):523, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Curnow:1978:SWS

[186] R. N. Curnow. Selection within self-fertilizing populations. Biometrics, 34 (4):603–610, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1978:GMC

[187] Anonymous. Gertrude Mary Cox — 1900–1978. Biometrics, 34(4):719– 720, 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1979:SR

[188] William G. Cochran. Some reflections. Biometrics, 35(1):1–2, 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:1979:GMC

[189] R. L. Anderson, R. J. Monroe, and L. A. Nelson. Gertrude M. Cox — a modern pioneer in statistics. Biometrics, 35(1):3–7, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prentice:1979:HRM

[190] Ross L. Prentice and J. D. Kalbfleisch. Hazard rate models with covariates. Biometrics, 35(1):25–39, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boyle:1979:MGM

[191] C. R. Boyle and Robert C. Elston. Multifactorial genetic models for quantitative traits in humans. Biometrics, 35(1):55–68, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lachenbruch:1979:DA

[192] P. A. Lachenbruch and M. Goldstein. Discriminant analysis. Biometrics, 35(1):69–85, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pollak:1979:SMG

[193] Edward Pollak. Some models of genetic selection. Biometrics, 35(1): 119–137, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Engen:1979:SAM

[194] Steinar Engen. Stochastic abundance models in ecology. Biometrics, 35 (1):331–338, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Purcell:1979:ESD

[195] Noel J. Purcell and Leslie Kish. Estimation for small domains. Biometrics, 35(2):365–384, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gail:1979:COL

[196] M. H. Gail and J. H. Ware. Comparing observed life table data with a known survival curve in the presence of random censorship. Biometrics, 35 (2):385–391, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hollander:1979:TDU

[197] Myles Hollander and Frank Proschan. Testing to determine the underlying distribution using randomly censored data. Biometrics, 35(2):393–401, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [215].

Andrich:1979:MCT

[198] David Andrich. A model for contingency tables having an ordered response classification. Biometrics, 35(2):403–415, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhapkar:1979:TMS

[199] Vasant P. Bhapkar. On tests of marginal symmetry and quasisymmetry in two- and three-dimensional contingency tables. Biometrics, 35(2): 417–426, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beck:1979:SSM

[200] Gerald J. Beck. Stochastic survival models with competing risks and covariates. Biometrics, 35(2):427–438, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:1979:AMN

[201] J. Rice, T. Reich, C. R. Cloninger, and R. Wette. An approximation to the multivariate normal integral: its application to multifactorial qualitative traits. Biometrics, 35(2):451–459, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:1979:STI

[202] B. Rosner, A. Donner, and C. H. Hennekens. Significance testing of interclass correlations from familial data. Biometrics, 35(2):461–471, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reed:1979:EDC

[203] George F. Reed and Richard B. McHugh. The estimation of dental caries incidence in the presence of diagnostic error. Biometrics, 35(2):473–478, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Knoke:1979:NAS

[204] James D. Knoke. Normal approximations for serial correlation statistics. Biometrics, 35(2):491–495, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shah:1979:MIL

[205] B. K. Shah. On the method of internal least squares. Biometrics, 35(2): 497–502, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burnham:1979:PGH

[206] K. P. Burnham. A parametric generalization of the Hayne estimator for line transect sampling. Biometrics, 35(3):587–595, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindley:1979:ALT

[207] D. V. Lindley. Analysis of life tables with grouping and withdrawals. Biometrics, 35(3):605–612, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Day:1979:THC

[208] N. E. Day and D. P. Byar. Testing hypotheses in case-control studies — equivalence of Mantel–Haenszel statistics and logit score tests. Biometrics, 35(3):623–630, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

North:1979:MHS

[209] Philip M. North and Byron J. T. Morgan. Modelling heron survival using weather data. Biometrics, 35(3):667–681, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1979:CIB

[210] William G. Cochran and David J. Finney. Chester Ittner Bliss (1899– 1979). Biometrics, 35(4):715–717, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1979:SBC

[211] J. S. Williams. A synthetic basis for a comprehensive factor-analysis theory. Biometrics, 35(4):719–733, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glaz:1979:PMA

[212] J. Glaz. Probabilities and moments for absorption in finite homogeneous birth-death processes. Biometrics, 35(4):813–816, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hauck:1979:LSV

[213] Walter W. Hauck. The large sample variance of the Mantel–Haenszel estimator of a common odds ratio. Biometrics, 35(4):817–819, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tavare:1979:NFH

[214] S. Tavaré. A note on finite homogeneous continuous-time Markov chains. Biometrics, 35(4):831–834, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hollander:1979:CTD

[215] M. Hollander and F. Proschan. Correction to: “Testing to determine the underlying distribution using randomly censored data” [Biometrics 35 (1979), no. 2, 393–401; MR 81a:62034]. Biometrics, 35(4):905, 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See